OpenStudy (anonymous):

can someone please help me evaluate lim 3xe^-x x->∞

geerky42 (geerky42):

Do you know L'Hopital's Rule?

OpenStudy (anonymous):

a bit. how would I start it though

geerky42 (geerky42):

Your limit is equal to \(\displaystyle \lim_{x\to\infty}\dfrac{3x}{e^x}\), right? Since direct substitution will get you \(\dfrac{\infty}{\infty}\) form, you can apply L'Hopital's Rule: \[\lim_{x\to\infty}\dfrac{3x}{e^x} = \lim_{x\to\infty}\dfrac{3}{e^x}\] Answer should be clear now.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay thanks :)